# Thunder-NUBench: A Benchmark for LLMs’ Sentence-Level Negation Understanding

### v1

@@ Abstract
Negation is a fundamental linguistic phenomenon that poses persistent challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Existing benchmarks often treat negation as a side case within broader tasks like natural language inference, resulting in a lack of benchmarks that exclusively target negation understanding. In this work, we introduce Thunder-NUBench, a novel benchmark explicitly designed to assess sentence-level negation understanding in LLMs. Thunder-NUBench goes beyond surface-level cue detection by contrasting standard negation with structurally diverse alternatives such as local negation, contradiction, and paraphrase. The benchmark consists of manually curated sentence-negation pairs and a multiple-choice dataset that enables in-depth evaluation of models’ negation understanding.
@@ 1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely linked to various human communicative abilities, including denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation is still prevalent in language; approximately 25% of sentences in English texts include some form of negation (Sarabi and Blanco, 2016; Hossain et al., 2020; Horn and Wansing, 2025).
Negation plays a crucial role in various natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI). Accurately interpreting negation is vital for understanding semantic oppositions (Khandelwal and Sawant, 2020; Hosseini et al., 2021; Singh et al., 2023). Recent research has shown that the importance of correctly handling negation extends even to multimodal language models (Quantmeyer et al., 2024; Alhamoud et al., 2025; Park et al., 2025), underscoring its widespread relevance across different domains.
@@
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find negated statements more difficult to process and understand compared to affirmative ones (Wales and Grieve, 1969; Sarabi and Blanco, 2016). Similarly, multiple studies have found that pretrained language models (PLMs) struggle to accurately interpret negation. For example, models like BERT (Devlin et al., 2019) and even large language models (LLMs) such as GPT-3 (Radford et al., ) frequently fail to differentiate between negated and affirmative statements. These models often rely on superficial cues, which can lead to incorrect outputs in the presence of negation (Kassner and Schütze, 2020; Hossain et al., 2022a; Truong et al., 2023).
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources treat negation as a minor aspect within broader tasks or focus solely on narrow syntactic detection. Consequently, evaluations have primarily been limited to encoder-based models (Hossain et al., 2020; Geiger et al., 2020; Truong et al., 2022; Anschütz et al., 2023).
To address these shortcomings, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset designed to assess large language models’ ability to interpret negation. The contributions of this paper are summarized as follows:
  • We define and categorize various negation phenomena, highlighting their differences from contradiction and paraphrase.
  • We introduce a manually curated benchmark to assess the ability of LLMs to understand these distinctions.
  • We perform systematic evaluations of several decoder-based LLMs using both prompting and fine-tuning approaches.
@@
Our benchmark offers valuable insights into the semantic reasoning abilities of language models and serves as a robust evaluation standard for future advancements in understanding negation.
@@ 2 Related Work
This section reviews existing studies on how language models understand and process negation.
@@
Early negation detection and scope resolution work focuses on rule-based systems and handcrafted heuristics, particularly in domain-specific settings like clinical texts. These systems are effective but lack flexibility across domains (Chapman et al., 2001; Carrillo de Albornoz et al., 2012; Ballesteros et al., 2012; Basile et al., 2012). Traditional machine learning methods, such as SVMs (Hearst et al., 1998) and CRFs (Sutton et al., 2012), are later introduced, though they also remain limited to narrow domains (Morante et al., 2008; Morante and Daelemans, 2009; Read et al., 2012; Li and Lu, 2018).
More recently, deep learning approaches employing CNNs (O’shea and Nash, 2015) and BiLSTM networks (Siami-Namini et al., 2019) have improved performance through better contextual embedding and sequence modeling (Fancellu et al., 2016; Bhatia et al., 2020). Pretrained transformer models like BERT (Devlin et al., 2019) have been leveraged through transfer learning (e.g., NegBERT (Khandelwal and Sawant, 2020)), significantly enhancing the accuracy of negation detection tasks. However, these methods still primarily address syntactic span detection, with deeper semantic comprehension of negation remaining challenging.
Negation-sensitive subtasks of NLU.
Negation understanding has become increasingly important in Natural Language Understanding (NLU) tasks (Hosseini et al., 2021). However, existing NLU benchmarks, such as SNLI (Bowman et al., 2015) for NLI, CommonsenseQA (Talmor et al., 2019) for QA, SST-2 (Socher et al., 2013) for sentiment analysis, STS-B (Cer et al., 2017) for textual similarity and paraphrasing, have been criticized for insufficiently accounting for the semantic impact of negation (Hossain et al., 2022a; Rezaei and Blanco, 2024). These datasets contain relatively few negation instances or include negations that are rarely critical to task performance, enabling language models to achieve high accuracy even when ignoring negation entirely.
Recent studies, such as NegNLI (Hossain et al., 2020), MoNLI (Geiger et al., 2020), NaN-NLI (Truong et al., 2022), have introduced negation-sensitive NLU benchmarks, demonstrating that model performance significantly declines when negation meaningfully affects the outcome (Naik et al., 2018; Yanaka et al., 2019; Hartmann et al., 2021; Hossain et al., 2022b; Hossain and Blanco, 2022; She et al., 2023; Anschütz et al., 2023). Such findings indicate that current language models heavily rely on superficial linguistic patterns rather than genuine semantic comprehension.
Limitations of distributional semantics.
Distributional semantics, the theoretical basis for many PLMs, is built on the distributional hypothesis. That is, words with similar meanings tend to occur in similar contexts (Harris, 1954; Sahlgren, 2008). This assumption enables models to learn semantic representations from textual co-occurrence patterns, making unsupervised training possible (Boleda, 2020; Lenci et al., 2022).
Although powerful in capturing broad semantic relationships, distributional semantics struggles significantly with negation because negated expressions (e.g., "not good") frequently occur in similar contexts as their affirmative forms ("good"), leading models to produce similar vector representations despite their opposite meanings. Previous studies have shown this limitation, highlighting how PLMs fail to capture semantic nuances introduced by antonyms and reversing polarity (Rimell et al., 2017; Jumelet and Hupkes, 2018; Niwa et al., 2021; Jang et al., 2022; Vahtola et al., 2022). Moreover, studies suggest that PLMs like BERT struggle to differentiate between affirmative and negated contexts (Kassner and Schütze, 2020; Ettinger, 2020).
@@
$\rightarrow$ He doesn’t speak English
        fluently.
@@
$\rightarrow$ He speaks English, but not
        fluently.

### v3

@@ Abstract
Negation is a fundamental linguistic phenomenon that poses ongoing challenges for Large Language Models (LLMs), particularly in tasks requiring deep semantic understanding. Current benchmarks often treat negation as a minor detail within broader tasks, such as natural language inference. Consequently, there is a lack of benchmarks specifically designed to evaluate comprehension of negation. In this work, we introduce Thunder-NUBench — a novel benchmark explicitly created to assess sentence-level understanding of negation in LLMs. Thunder-NUBench goes beyond merely identifying surface-level cues by contrasting standard negation with structurally diverse alternatives, such as local negation, contradiction, and paraphrase. This benchmark includes manually curated sentence-negation pairs and a multiple-choice dataset, allowing for a comprehensive evaluation of models’ understanding of negation.
@@ 1 Introduction
Negation is a fundamental and universal phenomenon found in languages worldwide. It is closely associated with various human communicative abilities, such as denial, contradiction, deception, misrepresentation, and irony. Although affirmative statements are more common, negation still plays a significant role in language; approximately 25% of sentences in English texts contain some form of negation (sarabi2016understanding; hossain2020analysis; sep-negation). This prevalence and its impact on meaning make accurate interpretation of negation crucial for several natural language processing (NLP) tasks, including sentiment analysis, question answering, knowledge base completion, and natural language inference (NLI) (khandelwal2020negbert; hosseini2021understanding; singh2023nlms). Recent studies have shown that effectively managing negation is important even for multimodal language models (quantmeyer2024and; alhamoud2025vision; park2025know).
@@
Meanwhile, negation poses significant challenges for both humans and language models. Research shows that people often find it more difficult to process and comprehend negated statements compared to affirmative ones (wales1969so; sarabi2016understanding). Similarly, many studies indicate that pretrained language models (PLMs) struggle to interpret negation accurately. For example, models like BERT (devlin2019bert) and even large language models (LLMs) such as GPT-3 (brown2020language) often have difficulty distinguishing between negated and affirmative statements. These models tend to rely on superficial cues, which can result in incorrect outputs when negation is involved (kassner2020negated; hossain2022analysis; truong2023language).
Figure 1: An example of Thunder-NUBench multiple-choice evaluation task, where the underlined text indicates the main verb phrase of each sentence, and the red text marks the negated part.
Despite its significance, there is a notable lack of dedicated evaluation benchmarks for understanding negation. Most existing resources either treat negation as a minor aspect of broader tasks or focus solely on narrow syntactic detection, often emphasizing encoder-based models (hossain2020analysis; geiger2020neural; truong2022not; anschutz2023not). To address this gap, we introduce Thunder-NUBench (Negation Understanding Benchmark), a dataset explicitly designed to evaluate LLMs’ sentence-level comprehension of negation. Our benchmark is structured as a multiple-choice question (MCQ) task: given an original sentence, the model must select the correct standard negation from four options. The other three choices—local negation, contradiction, and paraphrase—are carefully designed distractors that test whether models truly grasp semantic scope and logical oppositions.
The contributions of this paper are summarized as follows:
  • We define standard negation within the framework of sentential logic, moving beyond the cue-based and often ambiguous accounts of prior work. Grounding standard negation in logical structure not only clarifies its role in natural language but also supports the evaluation and enhancement of reasoning in LLMs.
  • We create a manually curated benchmark that includes a dataset of sentence-negation pairs for fine-tuning, along with a multiple-choice evaluation task.
  • We conduct systematic evaluations of decoder-based LLMs, assessing their performance under both prompting and supervised fine-tuning. This includes error and confusion analyses that highlight the models’ ongoing challenges with negation.
@@
Thunder-NUBench provides valuable insights into the semantic reasoning abilities of language models and serves as a robust standard for future research focused on understanding negation.
@@ 2 Related Work
Early work in negation detection and scope resolution primarily relied on rule-based systems and handcrafted heuristics, especially in domain-specific contexts like clinical texts. While these systems are effective, they lack flexibility across different domains (chapman2001simple; de2012ucm; ballesteros2012ucm; basile2012ugroningen). Traditional machine learning methods, such as Support Vector Machines (SVMs) (hearst1998support) and Conditional Random Fields (CRFs) (sutton2012introduction), were introduced later; however, they too are limited to narrow domains (morante2008learning; morante2009metalearning; read2012uio1; li2018learning).
More recently, deep learning approaches employing Convolutional Neural Networks (CNNs) (o2015introduction) and Bidirectional Long Short-Term Memory (BiLSTM) networks (siami2019performance) have enhanced performance by providing improved contextual embeddings and sequence modeling (fancellu2016neural; bhatia2019end). Pretrained transformer models like BERT have been employed through transfer learning techniques (e.g., NegBERT (khandelwal2020negbert)), significantly increasing the accuracy of negation detection tasks. Nonetheless, these methods still largely focus on syntactic span detection, leaving deeper semantic understanding of negation a challenging area to tackle.
@@
$\rightarrow$ He doesn’t speak English fluently.
@@
$\rightarrow$ He speaks English, but not fluently.